# Binarizing Physics-Inspired GNNs for Combinatorial Optimization

### v1

@@ 1 Introduction
With the advent of deep learning and its penetration into classical AI fields, such as automated planning and logical reasoning, combinatorial optimization (CO) has emerged as another promising area for its application. Among various deep learning approaches to CO, ranging from end-to-end machine learning algorithms to hybrid systems [11], physics-inspired graph neural networks (PI-GNNs) [47] have gained particular attention due to their generality and interpretable physical intuition. PI-GNNs leverage a specific loss function to encode the objective of an optimization problem, enabling unsupervised training to converge on solutions that minimize the objective. This formulation is highly adaptable across various combinatorial optimization domains, establishing PI-GNNs as a promising general-purpose CO-solving framework. The theoretical promise of PI-GNNs draws on insights from the overparameterization regime of neural networks, where multiple studies have shown that—under specific assumptions—the noise in stochastic gradient descent leads to convergence to global minima [60]. Coupled with their inherent parallelizability and computational power of modern GPUs, PI-GNNs emerge as a potentially more scalable and efficient alternative to classical solvers.
Despite their potential, PI-GNNs have recently faced critique for underperforming relative to efficient domain-specific heuristics on sparse graphs [1, 7]. In response, it has been argued that PI-GNNs exhibit superior performance on dense graph instances [49], considered better suited for the natively parallel deep learning-based approach [50]. However, the experimental analysis has focused merely on sparse $d$-regular graphs ($d=3$, $d=5$),111with edge density [15] $\rho(\mathcal{G})\leq 0.051$ across all experiments. Schuetz et al. [50] claim that $7$ MaxCut problems from the Gset dataset [59] are an exception, but their density is low, as well: $\rho(\mathcal{G})\leq 0.015$. Truly dense problems were tackled only within a different domain [48]. despite denser graphs having an essential status in domains involving network analysis [39, 42].
@@ 2 Background
We target physics-inspired (Sec. 2.2) GNN models (Sec. 2.3) for solving combinatorial optimization problems (Sec. 2.1).
@@ 2.1 Combinatorial Optimization
Figure 1: Schematic of an Ising spin chain in 1D, with arrow/spin down corresponding to assignment 00, and up to $1$, respectively.
@@
When Karp [28] introduced the hierarchy of $21$ combinatorial NP-complete problems, a significant subset shared a common structure, being defined over graphs. These include problems of maximum cut, maximum independent set, maximum clique, and graph partitioning, among others. Moreover, Karp demonstrated that even the remaining problems in this hierarchy can be polynomially reduced to one of these graph-based problems. Here we introduce two NP-hard CO problems that will be used in the paper.
@@ 2.2 Physics-Based Approaches to CO
A particularly popular spin-glass model relevant to combinatorial optimization is the Ising model. It describes a lattice of magnetic dipoles/spins (for a $1$D chain of Ising spins, see Fig. 1) that may take one of two discrete orientations. Its Hamiltonian can be written as:
@@
where $x_{i}\in\{\pm 1\}$ (or $\{0,1\}$, depending on conventions) are the spin variables and $J$ is a symmetric matrix of pairwise interaction coefficients. The second summand in the Ising Hamiltonian is the “magnetic field term” (with $h$ being a vector of local fields) that can incorporate penalty or constraint enforcement if mapped appropriately from an optimization problem, such as the MaxCut (Def. 2).
By encoding the cut as a spin arrangement, minimizing $H_{\rm Ising}$ aligns with finding an optimal partition. This has been the traditional way to approach CO problems with quantum computers, either with quantum annealing, as mentioned earlier, or with variational quantum algorithms such as the Quantum Approximate Optimization Algorithm (QAOA) [17]. The connecting characteristic of all the above is that the problem is first formulated as a quadratic unconstrained binary optimization (QUBO) [20] problem which has a natural mapping to the Ising model [36], as shown in Fig. 2. Such an approach has practical applications for industrial-scale problems [10, 18, 37] where classic methods, such as the B&B, are not sufficient.
@@ 2.3 Graph Neural Networks
The outlined close relationship between CO problems and graph representations (Sec. 2.1) highlights important characteristics that a solving approach should address. First, the problem representation should be permutation invariant, as the problem’s structure—defined by the graph’s connectivity—is essential, while the serialized order of its specification is irrelevant. Second, combinatorial optimization typically involves graphs of varying sizes and shapes. Therefore, a representation invariant to such structural changes is highly desirable.
@@
Solving CO problems with unsupervised learning generally received the most attention. Here, Karalias and Loukas [27] proposed a framework with probabilistic guarantees, which was later extended with principled objective relaxation [55] and meta-learning [54]. Orthogonally, the idea of PI-GNNs [47, 48] was developed, followed by multiple enhancements including reinforcement learning [45], automated machine learning [33], recurrent updates [44], quantum annealing [25, 34], efficient graph reduction [53], and generalization to mixed-integer linear programming [58].
@@
Figure 5: Schema of the PI-GNNs pipeline for solving CO under QUBO formalization (problem encoding, graph convolutions, node assignments), showcased on a MaxCut problem. The proposed modifications—binarization of the ultimate nonlinearity (Sec. 3.3) and loss fuzzification (Sec. 3.2)—are highlighted in blue.
@@ 3 The Approach
As outlined in Sec. 1, the performance of the PI-GNNs declines abruptly with the increasing density of the CO graphs, diminishing the promise of scaling to the more difficult problems through GPU parallelization. The contribution of this paper is to (i) provide an analysis of the issue (Sec. 3.1) and (ii) propose improvements (Sec. 3.2 and 3.3) to advance to the denser CO settings.
As a problem encoding, we consider the quadratic unconstrained binary optimization (QUBO; Sec. 2.2) model [20] which unifies a wide variety of NP-hard CO problems, forming the basis for the well-established PI-GNNs [47] framework we build upon in this work. The cost function for a QUBO problem can be expressed in compact form with the following Hamiltonian:
@@
Here, $x\in\{0,1\}^{N}$ is the vector of binary decision variables, and the QUBO matrix $Q$ is a square matrix that encodes the objective of the original problem [20]. Without loss of generality, the $Q$-matrix can be assumed to be symmetric or in upper triangular form. This Hamiltonian is famously related to the Ising spin model Hamiltonian (Eq. 3, Sec. 2.2). Its encoding (Eq. 5) for an example MaxCut problem (Def. 2) is displayed in Fig. 2.
With the PI-GNNs, an instance of a combinatorial optimization problem (Def. 1) is then given by a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with an associated adjacency matrix $A_{\mathcal{G}}$, and a cost function specified by the QUBO Hamiltonian $H_{Q}$. Input features (embeddings) of nodes $\mathcal{V}$ in $\mathcal{G}$ are randomly initialized, and fed together with $A_{\mathcal{G}}$ to a standard GNN model (Sec. 2.3) trained against the $H_{Q}$ loss minimization. During a single inference step, the graph is passed through multiple convolutional layers with intermediate ReLU nonlinearities. After the final graph layer, the (scalar) node outputs ($a^{pre}$) are projected into the $(0,1)$ interval using a logistic sigmoid $\sigma(a^{pre})$, serving as “soft node assignments” $a^{post}$ entering the loss $H_{Q}$. Once the training converges, these final output values $x_{i}:=a_{i}^{post}$ are thresholded as ${\rm int}(x_{i})$ to produce binary value assignments for the original CO problem.
@@ 3.1 Analyzing PI-GNNs Training Dynamics
To analyze the declining performance of PI-GNNs with increasing graph density, we tracked the model’s intermediate and final outputs during the training. Particularly, we tracked the “pre-activation” ($a^{pre}$) outputs of the final graph layer, and the “post-activation” ($a^{post}$) outputs of the final non-linearity, before the ultimate rounding (see Fig. 5). The evolution of activations ($a^{post}$) during training for one of the sparser settings is visualized as a heatmap, where color-encoded histograms represent the distribution of activation values over training epochs, [^2] Frequencies are computed from training runs on $20$ graphs. shown in Fig. 3 and in supplementary materials. The analysis indicates that, starting from random initial assignments, nodes (variables) first undergo a nullification phase during the initial thousands of iterations, minimizing the positive terms of the QUBO loss associated with constraint violations. Subsequently, a subset of nodes gradually becomes active in a later phase, increasing the negative terms of the QUBO loss that contribute to the solution value.
The later stage is, however, absent in the training dynamics on denser graphs ($d\in\{20,30,40,50\}$), resulting in a degenerated all-zero assignment, as shown in Fig. 4 (b). [^3] The difference in epoch range (cf. Fig. 3) is caused by early stopping. To inspect this failure mode more closely, a heatmap of pre-activation values from the same experiment is plotted in Fig. 4 (a).
Stagnation in training dynamics can typically be attributed to the presence of multimodality, that is, multiple wells of suboptimal local minima, as well as locally flat “saddle” neighborhoods. The greater the multimodality, the greater the chance of a suboptimal yet sufficiently steep well-trapping gradient flow. There is a natural correspondence to this observation in the literature on modeling combinatorial optimization as spin-glass energy (Sec. 2.2). Cheeseman et al. [12] observed that the vast majority of hard complexity combinatorial problems are, in practice, not especially problematic to solve. However, one can observe that certain perturbations of the combinatorial problem definition lead to starkly more difficult problems to solve, whose actual computational load for approaching a solution begins to approach the worst-case complexity. Simultaneously, along this perturbation, the corresponding energy of the spin glass model describing the combinatorial problem undergoes phase transitions, complicating the topology of the critical points of the objective landscape. We hypothesize that increasing graph connectivity, and thus the variable interdependencies, constitutes such a perturbation.
@@ 3.2 Fuzzy Logic Interpretation of the QUBO Relaxation
As outlined in Fig. 2, the cost function from Eq. 5 directly encodes graph optimization problems with binary-valued variables $x_{i}$. The interpretation is that the off-diagonal elements of the $Q$ matrix ($Q_{ij},i\neq j$) represent the penalties/rewards for edges included in the solution set, i.e., the penalty is added if both nodes are selected: $(x_{i}=1)\land(x_{j}=1)$. We will further use a simplified notation $x_{i}\land x_{j}$ or its algebraic version $x_{i}\cdot x_{j}$. The on-diagonal elements represent the penalties/rewards for nodes from the solution set, and can also be selected by a (trivial) $x_{i}\land x_{i}$.
In the relaxed PI-GNN version, the variables become real-valued $x_{i}\in[0,1]$, corresponding to “soft” assignments (node selections). To address the aforementioned problematic phase shift from 00 to $1$ (Sec. 3.1), we propose to inspect this binary relaxation in a more principled manner. Specifically, we propose to model the assignments as fuzzy values representing membership degrees $\mu(x_{i})$ of nodes to the solution set (Def. 1). With that interpretation, each term for the elements of matrix $Q$ now expresses a fuzzy conjunction.
@@
[…]
@@
[…]
The landscape contribution for each term in the first Hamiltonian will be a convex-bowed increasing function as $x_{i}$ and $x_{j}$ increase from 00 to $1$. For the second Hamiltonian, the loss is piecewise linear.
@@
Consider, by contrast, the fuzzy Lukasiewicz conjunction. This has two contrasting features. First, it is piecewise linear, and thus has no curvature, obviating the potential of producing saddle points. Second, for $x_{i}+x_{j}\leq 1$, which amounts to half of the measure of $[0,1]\times[0,1]$, the objective component is entirely flat at zero. This creates a structural sparsity in the dependence of the objective with respect to the inputs. Both of these factors mitigate some of the landscape complexity introduced by an increasing number of nonzero $\{Q_{ij}\}$ [14]. One can also observe that this activation resembles the topography of the piece-wise linear Rectified Linear Units. After [19], the use of these activations became more favored in NN models, observed to exhibit better performance and more stable training. See further Gombotz et al. [23], Mueller-Krumbhaar [41], Shafee [51] for applications of fuzzy conjunctions in spin glass models.
@@ 3.3 Binarizing PI-GNNs
Binarized neural networks (BNNs) [26] are a class of deep learning models with weights and activations constrained to binary values. Originally, BNNs have been constrained to values $1$ and $-1$ but, similarly to the QUBO and Ising models (Sec. 2.2), the methods can be adjusted to other pairs of values $(a,b)$, such as the logical $(0,1)$. The primary contribution of BNNs is that the transformations over purely binary matrices (implementable by bitwise operations) can reduce memory consumption and computational requirements [52]. However, importantly for this work, the binarized outputs also naturally represent discrete problems, avoiding the need for the relaxation to real values. Exploiting the latter advantage, we aim to binarize the output PI-GNNs activations, particularly the logistic sigmoid function $\sigma(x)=\frac{1}{1+e^{-x}}$ used by Schuetz et al. [47], resulting in a step function:
@@
Figure 6: Distributions of GNN’s pre-activation ($a^{pre}$) and post-activation ($a^{post}$) values over training epochs on $50$-regular MaxCut graphs: (a) $a^{pre}$ for bin-ste, (b) $a^{post}$ for bin-ste, (c) $a^{pre}$ for fuzzy-luk, (d) $a^{post}$ for fuzzy-luk.

### v2

@@ 1 Introduction
With the advent of deep learning and its penetration into classical AI fields, such as automated planning and logical reasoning, combinatorial optimization (CO) has emerged as another promising area for its application. Among various deep learning approaches to CO, ranging from end-to-end machine learning algorithms to hybrid systems [11], physics-inspired graph neural networks (PI-GNNs) [49] have gained particular attention due to their generality and interpretable physical intuition. PI-GNNs leverage a specific loss function to encode the objective of an optimization problem, enabling unsupervised training to converge on solutions that minimize the objective. This formulation is highly adaptable across various combinatorial optimization domains, establishing PI-GNNs as a promising general-purpose CO-solving framework. The theoretical promise of PI-GNNs draws on insights from the overparameterization regime of neural networks, where multiple studies have shown that—under specific assumptions—the noise in stochastic gradient descent leads to convergence to global minima [39, 62]. Coupled with their inherent parallelizability and computational power of modern GPUs, PI-GNNs emerge as a potentially more scalable and efficient alternative to classical solvers.
Despite their potential, PI-GNNs have recently faced critique for underperforming relative to efficient domain-specific heuristics on sparse graphs [1, 7]. In response, it has been argued that PI-GNNs exhibit superior performance on dense graph instances [51], considered better suited for the natively parallel deep learning-based approach [52]. However, the experimental analysis has focused merely on sparse $d$-regular graphs ($d=3$, $d=5$),111with edge density [15] $\rho(\mathcal{G})\leq 0.051$ across all experiments. Schuetz et al. [52] claim that 777 MaxCut problems from the Gset dataset [61] are an exception, but their density is low, as well: $\rho(\mathcal{G})\leq 0.015$. Truly dense problems were tackled only within a different domain [50]. despite denser graphs having an essential status in domains involving network analysis [41, 44].
@@ 2 Background
We target physics-inspired (Sec.˜2.2) GNN models (Sec.˜2.3) for solving combinatorial optimization problems (Sec.˜2.1).
@@ 2.1 Combinatorial Optimization
Figure 1: Schematic of an Ising spin chain in 1D, with arrow/spin down corresponding to assignment 0, and up to 111, respectively.
@@
When Karp [29] introduced the hierarchy of 212121 combinatorial NP-complete problems, a significant subset shared a common structure, being defined over graphs. These include problems of maximum cut, maximum independent set, maximum clique, and graph partitioning, among others. Moreover, Karp demonstrated that even the remaining problems in this hierarchy can be polynomially reduced to one of these graph-based problems. Here we introduce two NP-hard CO problems that will be used in the paper.
@@ 2.2 Physics-Based Approaches to CO
A particularly popular spin-glass model relevant to combinatorial optimization is the Ising model. It describes a lattice of magnetic dipoles/spins (for a  [^1] D chain of Ising spins, see Fig.˜1) that may take one of two discrete orientations. Its Hamiltonian can be written as:
@@
[…]
By encoding the cut as a spin arrangement, minimizing $H_{\rm Ising}$ aligns with finding an optimal partition. This has been the traditional way to approach CO problems with quantum computers, either with quantum annealing, as mentioned earlier, or with variational quantum algorithms such as the Quantum Approximate Optimization Algorithm (QAOA) [17]. The connecting characteristic of all the above is that the problem is first formulated as a quadratic unconstrained binary optimization (QUBO) [20] problem which has a natural mapping to the Ising model [37], as shown in Fig.˜2. Such an approach has practical applications for industrial-scale problems [10, 18, 38] where classic methods, such as the B&B, are not sufficient.
@@ 2.3 Graph Neural Networks
The outlined close relationship between CO problems and graph representations (Sec.˜2.1) highlights important characteristics that a solving approach should address. First, the problem representation should be permutation invariant, as the problem’s structure—defined by the graph’s connectivity—is essential, while the serialized order of its specification is irrelevant. Second, combinatorial optimization typically involves graphs of varying sizes and shapes. Therefore, a representation invariant to such structural changes is highly desirable.
@@
Solving CO problems with unsupervised learning generally received the most attention. Karalias and Loukas [28] proposed a framework with probabilistic guarantees, later extended with principled objective relaxation [57] and meta-learning [56]. Orthogonally, the idea of PI-GNNs [49, 50] was developed, followed by enhancements including reinforcement learning [47], automated machine learning [34], recurrent updates [46], quantum annealing [25, 35], efficient graph reduction [55], and generalization to mixed-integer linear programming [60]. Finally, Ichikawa [27] focused on annealing the continuous relaxation via a loss term.
@@
Figure 5: Schema of the PI-GNNs pipeline for solving CO under QUBO formalization (problem encoding, graph convolutions, node assignments), showcased on a MaxCut problem. The proposed modifications—binarization of the ultimate nonlinearity (Sec.˜3.3) and loss fuzzification (Sec.˜3.2)—are highlighted in blue.
@@ 3 The Approach
As outlined in Sec.˜1, the performance of the PI-GNNs declines abruptly with the increasing density of the CO graphs, diminishing the promise of scaling to the more difficult problems through GPU parallelization. The contribution of this paper is to (i) provide an analysis of the issue (Sec.˜3.1) and (ii) propose improvements (Sec.˜3.2 and 3.3) to advance to the denser CO settings.
As a problem encoding, we consider the quadratic unconstrained binary optimization (QUBO; Sec.˜2.2) model [20] which unifies a variety of NP-hard CO problems, forming the basis for the PI-GNNs [49] framework we build upon in this work. The cost function for a QUBO problem can be expressed in compact form with the Hamiltonian:
@@
Here, $x\in\{0,1\}^{N}$ is the vector of binary decision variables, and the QUBO matrix $Q$ is a square matrix that encodes the objective of the original problem [20]. Without loss of generality, the $Q$-matrix can be assumed to be symmetric or in upper triangular form. This Hamiltonian is famously related to the Ising spin model Hamiltonian (Eq.˜3, Sec.˜2.2). Its encoding (Eq.˜5) for an example MaxCut problem (Def.˜2) is displayed in Fig.˜2.
With the PI-GNNs, an instance of a combinatorial optimization problem (Def.˜1) is then given by a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with an associated adjacency matrix $A_{\mathcal{G}}$, and a cost function specified by the QUBO Hamiltonian $H_{Q}$. Input features (embeddings) of nodes $\mathcal{V}$ in $\mathcal{G}$ are randomly initialized, and fed together with $A_{\mathcal{G}}$ to a standard GNN model (Sec.˜2.3) trained against the $H_{Q}$ loss minimization. During a single inference step, the graph is passed through multiple convolutional layers with intermediate ReLU nonlinearities. After the final graph layer, the (scalar) node outputs ($a^{pre}$) are projected into the (0,1)(0,1)( 0 , 1 ) interval using a logistic sigmoid $\sigma(a^{pre})$, serving as “soft node assignments” $a^{post}$ entering the loss $H_{Q}$. Once the training converges, these final output values $x_{i}:=a_{i}^{post}$ are thresholded as ${\rm int}(x_{i})$ to produce binary value assignments for the original CO problem.
@@ 3.1 Analyzing PI-GNNs Training Dynamics
To analyze the declining performance of PI-GNNs with increasing graph density, we tracked the model’s intermediate and final outputs during the training. Particularly, we tracked the “pre-activation” ($a^{pre}$) outputs of the final graph layer, and the “post-activation” ($a^{post}$) outputs of the final non-linearity, before the ultimate rounding (see Fig.˜5). The evolution of activations ($a^{post}$) during training for one of the sparser settings is visualized as a heatmap, where color-encoded histograms represent the distribution of activation values over training epochs, [^2] Frequencies are computed from training runs on 202020 graphs. shown in Fig.˜3 and in App.˜G. The analysis indicates that, starting from random initial assignments, nodes (variables) first undergo a nullification phase during the initial thousands of iterations, minimizing the positive terms of the QUBO loss associated with constraint violations. Subsequently, a subset of nodes gradually becomes active in a later phase, increasing the negative terms of the QUBO loss that contribute to the solution value.
The later stage is, however, absent in the training dynamics on denser graphs ($d\in\{20,30,40,50\}$), resulting in a degenerated all-zero assignment, as shown in Fig.˜4 (b). [^3] The difference in epoch range (cf. Fig. 3) is caused by early stopping. To inspect this failure mode more closely, a heatmap of pre-activation values from the same experiment is plotted in Fig.˜4 (a).
Stagnation in training dynamics can typically be attributed to the presence of multimodality, that is, multiple wells of suboptimal local minima, as well as locally flat “saddle” neighborhoods. The greater the multimodality, the greater the chance of a suboptimal yet sufficiently steep well-trapping gradient flow. There is a natural correspondence to this observation in the literature on modeling combinatorial optimization as spin-glass energy (Sec.˜2.2). Cheeseman et al. [12] observed that the vast majority of hard complexity combinatorial problems are, in practice, not especially problematic to solve. However, one can observe that certain perturbations of the combinatorial problem definition lead to starkly more difficult problems to solve, whose actual computational load for approaching a solution begins to approach the worst-case complexity. Simultaneously, along this perturbation, the corresponding energy of the spin glass model describing the combinatorial problem undergoes phase transitions, complicating the topology of the critical points of the objective landscape. We hypothesize that increasing graph connectivity, and thus the variable interdependencies, constitutes such a perturbation.
@@ 3.2 Fuzzy Logic Interpretation of the QUBO Relaxation
As outlined in Fig.˜2, the cost function from Eq.˜5 directly encodes graph optimization problems with binary-valued variables $x_{i}$. The interpretation is that the off-diagonal elements of the $Q$ matrix ($Q_{ij},i\neq j$) represent the penalties/rewards for edges included in the solution set, i.e., the penalty is added if both nodes are selected: $(x_{i}=1)\land(x_{j}=1)$. We will further use a simplified notation $x_{i}\land x_{j}$ or its algebraic version $x_{i}\cdot x_{j}$. The on-diagonal elements represent the penalties/rewards for nodes from the solution set, and can also be selected by a (trivial) $x_{i}\land x_{i}$.
In the relaxed PI-GNN version, the variables become real-valued $x_{i}\in[0,1]$, corresponding to “soft” assignments (node selections). To address the aforementioned problematic phase shift from 0 to 111 (Sec.˜3.1), we propose to inspect this binary relaxation in a more principled manner. Specifically, we propose to model the assignments as fuzzy values representing membership degrees $\mu(x_{i})$ of nodes to the solution set (Def.˜1). With that interpretation, each term for the elements of matrix $Q$ now expresses a fuzzy conjunction.
@@
Schuetz et al. [49] intuitively defaulted to the algebraic version of the classical conjunction $x_{i}\cdot x_{j}$, corresponding to a product fuzzy conjunction $x_{i}\mathbin{\ooalign{$\land$\cr\kern 2.15277pt\cr$\scriptscriptstyle P\mkern 1.5mu$\cr}}x_{j}$. However, there is a number of other common $t$-norms, such as the standard $x_{i}\mathbin{\ooalign{$\land$\cr\kern 2.15277pt\cr$\scriptscriptstyle S\mkern 1.5mu$\cr}}x_{j}=\min\{x_{i},x_{j}\}$ or Łukasiewicz conjunction $x_{i}\mathbin{\ooalign{$\land$\cr\kern 2.15277pt\cr$\scriptscriptstyle L\mkern 1.5mu$\cr}}x_{j}=\max\{x_{i}+x_{j}-1,0\}$, which can be more appropriate in different scenarios, including QUBO. The generalized fuzzy form of the Hamiltonian then reads:
@@
The landscape contribution for each term in the first Hamiltonian will be a convex-bowed increasing function as $x_{i}$ and $x_{j}$ increase from 0 to 111. For the second Hamiltonian, the loss is piecewise linear.
@@
Consider, by contrast, the fuzzy Lukasiewicz conjunction. This has two contrasting features. First, it is piecewise linear, and thus has no curvature, obviating the potential of producing saddle points. Second, for $x_{i}+x_{j}\leq 1$, which amounts to half of the measure of $\times[0,1]$ [ 0 , 1 ], the objective component is entirely flat at zero. This creates a structural sparsity in the dependence of the objective with respect to the inputs. Both of these factors mitigate some of the landscape complexity introduced by an increasing number of nonzero $\{Q_{ij}\}$ [14]. One can also observe that this activation resembles the topography of the piece-wise linear Rectified Linear Units. After [19], the use of these activations became more favored in NN models, observed to exhibit better performance and more stable training. See further Gombotz et al. [23], Mueller-Krumbhaar [43], Shafee [53] for applications of fuzzy conjunctions in spin glass models.
@@ 3.3 Binarizing PI-GNNs
Binarized neural networks (BNNs) [26] are a class of deep learning models with weights and activations constrained to binary values. Originally, BNNs have been constrained to values 111 and 1-1- 1 but, similarly to the QUBO and Ising models (Sec.˜2.2), the methods can be adjusted to other pairs of values $(a,b)$, such as the logical (0,1)(0,1)( 0 , 1 ). The primary contribution of BNNs is that the transformations over purely binary matrices (implementable by bitwise operations) can reduce memory consumption and computational requirements [54]. However, importantly for this work, the binarized outputs also naturally represent discrete problems, avoiding the need for the relaxation to real values. Exploiting the latter advantage, we aim to binarize the output PI-GNNs activations, particularly the logistic sigmoid function $\sigma(x)=\frac{1}{1+e^{-x}}$ used by Schuetz et al. [49], resulting in a step function:
@@
Figure 6: Distributions of GNN’s pre-activation ($a^{pre}$) and post-activation ($a^{post}$) values over training epochs on 505050-regular MaxCut graphs: (a) $a^{pre}$ for bin-ste, (b) $a^{post}$ for bin-ste, (c) $a^{pre}$ for fuzzy-luk, (d) $a^{post}$ for fuzzy-luk.